MonJoliNom MaJolieAdresse MonJoliTéléphone

June 22, 1999

LeJoliDestinataire SaJolieAdresse SonJoliTéléphone

Cher JoliDestinataire

Le corps de la lettre avec des jolies phrases.

MesJoliesSalutations

MonJoliNom

cc: UnJoliAutreDestinataire EncoreUnJoliDestinataire

PS: MonJoliPostScriptum